# Improving Expressivity of Graph Neural Networks using Localization

### v3

@@ 6.4 Algorithms for subgraph counting
Triangle counting in the host graph
We describe an algorithm for counting the number of triangles in a given host graph $G$ in Algorithm 5. Note that counting the number of triangles is the same as counting the number of edges in the subgraph induced by $N_{G}(v).$ It is well known that two $1-$WL equivalent graphs have the same number of edges. This ensures that if we run $1-$WL on the induced subgraphs in the neighborhood of $v$, taking color as a feature, we can guarantee the count of the triangles. On the other hand, we can see that running $1$- WL on graph $G$ will not guarantee the count of triangles. Running $1-$WL on the entire graph takes $O(n+m)\log(n)$ and $O(n)$ space, where $m$ is the number of edges. Thus, running $1-$WL locally in the neighborhood is more space and time efficient. Note that the running time is also dependent on the number of iterations, $h.$ Running $1-$WL for $h-$ iteration requires $O(nh)$ time. The quantifier depth of counting logic with $(k+1)$ variables is equivalent to the number of iterations of $k-$WL ( See Theorem 2). For the case of triangle counting, we just need to count the number of edges, which can be done by running just one iteration of $1-$WL. So, the time required is $O(deg(v))$ for each $v.$ This can be done in parallel for each vertex.
@@
Counting subgraph of radius one
We begin by explaining the procedure for counting the number of subgraphs having a dominating vertex (radius one). For this purpose, we fix a dominating vertex $u$. If a subgraph exists, then the dominating vertex must be mapped to some vertex. We iteratively map the dominating vertex to each vertex in the host graph and count the number of patterns in the neighborhood of the dominating vertex.
Here we present an algorithm for counting patterns of radius one Algorithm 6. Note that running $k-$WL on the entire graph takes $O(k^{2}\cdot n^{k+1}\log n)$ time and $O(n^{k})$ space, whereas when we run locally, it requires less time and space. Suppose we run only on the neighborhood of each vertex. Then, it requires $\sum_{v\in V(G)}(deg(v))^{k+1}\log(deg(v))$ and space $O(max_{i}(deg(v_{i}))^{k}+n).$ More specifically, suppose the given graph is $r$-regular. Then it requires $O(r^{k+1}\log(r)n)$ time and $O(r^{k}+n)$ space. Therefore, if the graph is sparse, then we can implement Local $k-$WL for a larger value of $k$. We can see that running $k-$WL locally is not dependent on the size of the graph exponentially. However, it is not feasible to run $k-$WL on the entire graph for a larger value of $k.$
@@
Counting subgraphs of radius r
Here, in Algorithm 7, we describe how to count the number of subgraphs of radius $r.$ We iterate over all the vertices and take the $r$-hop neighborhood around that vertex, say $v,$ and choose a suitable $k$ according to the structure of the pattern that can guarantee the count of subgraphs in the local substructure.
@@
6:         if ##\## vertices at layer $i$ from $v$ $<$ ##\## vertices at layer $i$ from $u$ then
@@
7 Fragmentation
Here, we discuss the Fragmentation technique that is different from the localized methods we have seen so far. From Table 1, we have seen that Local $k-$WL (or ((Local + Layer) $k-$WL)) is sufficient for getting an exact count for the patterns given in the table. Given a pattern $P,$ that is more complicated than the patterns in Table 1, we fragment $P$ into simpler patterns such that their exact count is known. In the subgraph GNN proposed earlier, look into subgraph of the host graph. We have seen that this technique is scalable on large graphs. Also, we have seen that subgraph GNN is more expressive and efficient than traditional GNN. So, we tried to explore the expressibility when the pattern is also fragmented into smaller subpatterns.
The fragmentation method involves fragmenting the pattern, $P$, into smaller subpatterns and counting these subpatterns to get the count of the $P$ in the host graph. As described in Section 6, the value of $k$ depends on the size of the pattern (induced subgraph) and its structure (subgraph count). Thus, even though the $htw(P)$ may be large, if we divide it into subpatterns, then $k$ is required to guarantee the count would be reduced. Thus, it provides more expressiveness for smaller $k$ in order to count the patterns that cannot be counted if we directly apply Local $k-$WL. Thus, given a pattern, we apply the same fragmentation on $G_{v}^{r}.$ Thus, the number of occurrences of $H$ in $G_{v}^{r}$ can be computed by combining the counts of the simpler patterns. Instead of training a GNN for counting $H,$ we can design GNNs for learning the easier tasks (i.e., for counting the simpler models) and combine the outputs of those models. It should be noted that the fragmentation into smaller subgraphs depends on the structure of the pattern $H.$
We demonstrate this technique for counting induced tailed triangles in Figure 2. As seen in the figure, the tailed triangle pattern can be fragmented into two parts : the pendant vertex as the key vertex, and an attributed triangle. The colors assigned to the nodes of the attributed triangle depend on the distance of the nodes from the key node. Thus, the task of counting tailed triangles reduces to counting attributed triangles, as all the vertices at level 1 are connected to the root.
Suppose the task is to count the number of chordal cycles appearing as induced subgraphs. If we pick the vertex of degree three as the key vertex, then it is enough to search the neighbourhood of $v$ in the host graph. Now, in $N_{G}(v),$ if we count the number of $2-stars,$ then it gives the count of chordal cycle appearing as subgraph. If we eliminate the appearance of $K_{4}$, then it would give the exact count of chordal cycles appearing as induced subgraphs. In that case, we count the number of triangles in $N_{G}(v),$ which gives the exact count of $K_{4}$.
Using the fragmentation technique, we show that just $1-$WL is sufficient for getting exact counts of induced subgraphs of certain sizes.
Figure 2: Fragmentation for counting tailed triangles.
Idea of the fragmentation algorithm 8: Given a graph $G$ and a pattern $P$, we first fix a vertex $u\in V(P)$ as the key vertex. Now, assume that the radius of the pattern is $r$. Thus, for counting $P$ locally, it is sufficient to take the $r$-hop neighbourhood for each vertex $v$ of $G,$ say $G_{v}^{r},$ as has been shown in Lemma 3 and Lemma 5. Also, we have proved above that doing count locally is sufficient for both subgraph and induced subgraph. Now, we fragment pattern $P$ into smaller subpatterns, say $P_{1},P_{2},\ldots P_{l}$. Based on the structure of $P$, we consider the subgraphs of $G_{v}^{r}$ where the subpattern $P_{i}$ is required to be counted. For all subpattern $P_{i}$ in $P$, we make a list of subgraphs $G_{v}^{r}(1),G_{v}^{r}(2),\ldots G_{v}^{r}(t)$ of $G_{v}^{r}$ where $P_{i}$ needs to be counted. We aggregate these lists into a dictionary, $\mathcal{L},$ with $P_{i}$ as the keys. It should be noted that the decomposition of $P$ into $P_{i}$s is such that their counts can be calculated. That is, we have learnt models $M_{i},$ corresponding to each $P_{i}$, which counts the number of subpatterns $P_{i}$. The array $c$ stores the count of $P_{i}$’s in each subgraph of $G_{v}^{r}$. Now, for each vertex, we use appropriate functions $\alpha$ and $\beta$ to combine the counts in $c$ to get the count of $P$ in $G_{v}^{r}$. Finally, the function $\gamma$ finds the normalizing factor to get the actual count of the pattern in the $G.$
@@ 7 Fragmentation
1:Let $G$ be the host graph and $P$ be the list of patterns,
$\mathcal{L}$ be the dictionary of subgraph rules associated with the subpattern $P_{i},$
3:$M=\{M_{1},\ldots,M_{l}\}$ the list of learnt models for counting $P_{i}$’s, where $l=|P|$
4:$a\leftarrow$ Zero array of size $|V(G)|$
5:for each vertex $v$ in $V(G)$ do
6:     Extract $G_{v}^{r}$
7:     $b\leftarrow$ Zero array of size $l$
8:     for each pattern $P_{i}$ in $\mathcal{L}$ do
9:         $c\leftarrow$ Zero array of size $s$, where $s=|\mathcal{L}(P_{i})|$
10:         for each rule $k$ in $\mathcal{L}(P_{i})$ do
11:              Extract $G_{v}^{r}(k)$
12:              $c[k]$ = $M_{i}(G_{v}^{r}(k))$
@@
14:         $b[i]$ = $\alpha(c)$
@@
16:     $a[v]$ = $\beta(b)$
17:end for
18:$Count(P,G)$ = $\gamma(a)$
19:return $Count(P,G)$
@@
Using the fragmentation method, we can count induced tailed triangle, chordal $C_{4}$ and $3-star,$ by running Local $1-$WL.
@@
For tailed triangle, we fix the pendant vertex as the key vertex (refer to Figure 2). Now, we have to look for the count of triangles such that exactly one of the node of triangle is adjacent to the key vertex. We find induced subgraph of $2$-hop neighborhood for each key vertex. Now, we color the vertices at distance $i$ by color $i$. Then, the problem of counting the number of tailed triangles reduces to counting the number of colored triangles in the colored graph such that one node of triangle is colored $1$ and remaining two nodes are colored $2$. We can find the count of colored triangles using $1-$WL on the induced subgraph by Corollary 1. This number of count of colored triangle is same as $IndCount_{(u,v}(tailed-triangle,G_{v}^{r})$. Now, using Lemma 3, we can say that fragmentation technique can count tailed triangles appearing as induced subgraphs using $1-$WL.
Consider the pattern, chordal $C_{4}.$ We have proved that $1-$WL can count the number of subgraphs of chordal $C_{4}.$ So, to count the number of induced subgraphs of chordal $C_{4},$ we only have to eliminate the count of $K_{4}.$ When we fix one vertex of degree $3$ as key vertex, we can easily compute the count of $K_{1,2}$ in the neighborhood. Now, we have to eliminate all three tuples appearing as triangles in the neighborhood of the key vertex. We can easily count the number of triangles in the neighborhood of each vertex. This gives the exact count of chordal $C_{4}$ appearing as subgraph in the local structure. Using Lemma 3, we can find $IndCount(ChordalC_{4},G)$.
Consider the pattern, $3-star$. Here, we choose any of the pendant vertex as key vertex. Now, we have to compute $K_{1,2},$ where one center vertex of the star is connected to the key vertex. We can easily count the number of colored $K_{1,2}$ in $2$-hop neighborhood of the key vertex. However, a triangle can be also be included in this count. So, we have to eliminate the $3$ tuples forming a triangle. Again, using the approach discussed above, we can count the number of colored triangles and this will output the exact count of colored induced $K_{1,2}$. Again, using lemma 3, we can find $IndCount(3-star,G).$
Theorem 11.
Proof.
In Table 2, we describe how the fragmentation technique can be leveraged to count all the induced subgraphs of size $4$. This shows that for $k=1$, the fragmentation technique is more expressive than $S_{k+3}$. Also, we can enlist more graphs where the larger pattern can be seen as the union of fragments of smaller patterns. Using this, we can see that it covers all the graphs that were mentioned in Chen et al. (2020). One can see that all the formulae use the function that can be computed by $1-$WL. The number of vertices and the number of edges can easily be computed after doing one iteration of $1-$WL. Also, the degree sequence can be computed after running $1-$WL for one iteration. All other functions(formulae) are just the linear combination of the functions computed earlier, the number of vertices or the number of edges. ∎
In the structures shown in the table 2 below, we have highlighted the key vertex by "light green" color and other vertices by "black color".
Table 2: Fragmentation for patterns for size of at most 4.
@@
G3 [Uncaptioned image] $\large\frac{\sum_{v}IndCount_{(u,v)}(G_{1},G-G[N_{G}[v]])}{3}$
@@ 8 Theoretical Comparison of Graph Neural Networks
In this section we give a comparison of the time and expressiveness between the GNN hierarchies proposed in Papp and Wattenhofer (2022a) and our methods. From Theorem 3, it is clear that $k-$WL is less expressive than Local $k-$WL. Also, we have shown that the space and time required by Local $k-$WL are less compared to $k-$WL.
@@
$S_{k}$ is a model in which a vertex is given an attribute based on the number of connected induced subgraphs of size at most $k,$ the key vertex. Even though it searches locally, the number of non-isomorphic graphs may be too many for a small value of $k$. Suppose the radius of the induced subgraph is $r$; then it has to search up to the $r$-hop neighborhood. Using brute force would require $O(n_{1}^{k})$-time to count the number of induced subgraphs of size $k,$ for each individual induced subgraph. To improve the time complexity, it either stores the previous computation, which requires a lot of space or further recomputes from scratch. Thus, it would require $O(t_{k}\times n_{1}^{k})$, where $t_{k}$ is the number of distinct graphs upto isomorphism of at most $k$-vertices. Using Theorem 6, one can easily see that running $k-$WL locally is more expressive than $S_{k+2}.$
@@
The $N_{k}$ model has a preprocessing step in which it takes the $k$-hop neighborhood around vertex $v,$ and gives attributes based on the isomorphism type. In a dense graph, it is not feasible to solve the isomorphism problem in general, as the size of the induced subgraph may be some function of $n$. Also, using the best known algorithm for graph isomorphism by Babai (2016), the time required is $O(n_{1}^{O(\log n_{1})}).$ However, running Local $k-$WL would require $O(n_{1}^{k})$. Also, there are rare examples of graphs that are $3-$WL equivalent and non-isomorphic. So, if we run $3-$WL locally, then most of times expressive power matches with $N_{k}.$
@@
The $M_{k}$ model deletes a vertex $v$ and then runs $1-WL.$ Papp and Wattenhofer (2022a) proposed that instead of deleting the vertices, $k$ vertices are marked in the local neighborhood and showed that it is more expressive than deletion. It identifies $k$ set of vertices in the local $r$-hop neighborhood of the graph. It would require $O(n_{1}^{(k+2)}\log(n_{1}))$ time as it has $O(n_{1}^{k})$ many possibilities of choosing $k$ many vertices. It requires $O(n^{2}\log n)$ time to run $1-$WL. The same time is required for Local $(k+1)-$WL. That’s why we compare $M_{k-1}$ with Local $k-$WL in Table 3. Also, it is known that with any $l$ markings and running $k-$WL is less expressive than running $(k+l)-$WL on the graph Fürer (2017). So, if we plug in the value, we can see that running Local $k-$WL is more expressive than doing $l$ marking and running $1-$WL. One can get an intuition by comparing with the $(k+1)$ bijective pebble game. If we mark the vertices, then the marking pebbles get fixed and give less power to spoiler. However, just running $k-$WL the spoiler is free to move all the pebbles. We present a simple proof that Local $k-$WL is at least as expressive as $M_{k-1}$.
@@
Let $G_{v}^{r}$ and $G_{u}^{r}$ be the induced subgraphs around the $r$-hop neighborhood for vertices $v$ and $u$, respectively. Let $M_{k}$ distinguish $G_{v}^{r}$ and $G_{u}^{r}$. We claim that Local $k-$WL can also distinguish the graph $G_{v}^{r}$ and $G_{u}^{r}$. To prove our claim, we use the pebble bijective game. $G_{v}^{r}$ is distinguished because there exists a tuple $(v_{1},v_{2},....v_{k-1})$ such that marking these vertices and running $1-$WL on the graph gives stabilised coloring of the vertices that does not match with that of $G_{u}^{r}.$ Now, consider two games. One game corresponds to $1-$WL and another to Local $k-$WL. For the first $(k-1)$ moves, the Spoiler chooses to play and places pebbles at $(v_{1},v_{2},....v_{k-1}).$ After that, in both games, there are two pebbles and the position of both the games are same. Let $S_{1}$ and $D_{1}$ be the spoiler and Duplicator in the $(k+1)$ bijective pebble game, and $S_{2}$ and $D_{2}$ be the spoiler and Duplicator in $2$ bijective pebble game. $S_{1}$ will follow the strategy of $S_{2}$ and $D_{2}$ follows the strategy of $D_{1}.$ We prove by induction on the number of rounds. Our induction hypothesis is that the position of games in both the games is the same and if $S_{2}$ wins, then $S_{1}$ also wins.
@@
Table 3: Here, $n:$ number of nodes in the graph, $n_{1}:$ max number of nodes in a $r$-hop neighbourhood of a vertex, $n_{2}:$ maximum number of nodes in any two consecutive layers for a particular vertex, over all the vertices of the graph, and $t_{k}:$ number of distinct graphs upto isomorphism of at most $k$-vertices. The first row compares the expressiveness of the models , and the second row compares the time complexity, with respect t0 our models.
Local k -WL (Local+Layer) k -WL k- WL S_k M_k-1
Expressiveness - - Less
Less than
Local $(k-2)-$WL
Less
Time $O(n\times n^{k+1}_{1})$ $O((n_{2})^{k+1}\times rn)$ $O(n^{k+1})$ $O(t_{k}n^{k}_{1}n)$ $O(n^{k+1}_{1}n)$
9 Proposed Architecture
In this section, we propose two GNN architectures for subgraph as well as induced subgraph counting. The design of these networks follows from the discussions about subgraph and induced subgraph counting in the previous sections. Given a graph $G$ as an input, we extract the rooted subgraphs $G_{v}^{r}$ for each vertex $v.$ Each of these subgraphs is rooted at $v.$ The value of $r$ depends on the structure to be counted. For example, we extract $1-$hop induced subgraphs for counting triangles. These subgraphs are then given as inputs to the proposed models.
9.1 InsideOut GNN
We conducted the experiments using two different architectures, InsideOutGNN and InSigGNN.
Figure 3: Schematic of the InsideOutGNN Model
The InsideOut GNN model (Figure 3) consists of an internal network and an external network. The internal network learns the local substructure counts for each rooted subgraph, $G_{v}^{r}.$ A subgraph are given as input pass through an Internal GNN model. This model consists of GINConv Xu et al. (2019) layers. The Internal GNN outputs the embeddings of the nodes present in a given subgraph. These node embeddings are then aggregated using a GlobalAddPool layer to get the embedding of the root node $v$ of the subgraph. This root node embedding is the feature vector for $v,$ as given in line $7$ of Algorithm 1. The GINConv layers simulate the WL iterations on the nodes of the subgraph. This root node embedding, $emb_{v},$ is then passed through a linear transformation layer to learn the local counts of the required substructure in the subgraph.
Now, we take a union of all the subgraphs for a given graph. Using the node embeddings, $emb_{v},$ learnt from the Internal GNN model, the External GNN model learns the global count of the required substructure. The External GNN model also comprises of GINConv layers.
The intuition behind this design is to make the models learn the local counts well. Once these local counts are learned, the global count is just a linear function of the local counts, which can be learnt easily by the External GNN model.
9.2 InSigGNN
For the InSig model, the Internal GNN component is similar to that of the InsideOut GNN model. The architecture is shown in Figure 4. However, in this model, we do not transfer the node embeddings learnt from the Internal GNN part of the model. Instead, we add all the local counts predicted by the internal part of the model and pass the sum as an input to the external model, which is just a linear transformation layer. The weights learned in the Internal GNN model depend only on the local counts of the substructures in the rooted subgraphs. The external linear transformation is learned based on the global count of a substructure in the entire input graph.
Figure 4: Schematic of the InSigGNN model

### v5

@@ 6.4 Algorithms for subgraph counting
6.4.1 Triangle counting in the host graph
We describe an algorithm for counting the number of triangles in a given host graph $G$ in 5. Note that counting the number of triangles is the same as counting the number of edges in the subgraph induced by $N_{G}(v).$ It is well known that two 11-WL equivalent graphs have the same number of edges. This ensures that if we run $1$-WL on the induced subgraphs in the neighbourhood of $v$, taking colour as a feature, we can guarantee the count of the triangles. On the other hand, we can see that running 11- WL on graph $G$ will not guarantee the count of triangles. Running 11-WL on the entire graph takes $O(n+m)\log(n)$ and $O(n)$ space, where $m$ is the number of edges. Thus, running $1$-WL locally in the neighbourhood is more space and time-efficient. Note that the running time is also dependent on the number of iterations, $h.$ Running 11-WL for $h$- iteration requires $O(nh)$ time. The quantifier depth of counting logic with $(k+1)$ variables is equivalent to the number of iterations of $k$-WL ( See 2). For the case of triangle counting, we just need to count the number of edges, which can be done by running just one iteration of 11-WL. So, the time required is $O(deg(v))$ for each $v.$ This can be done in parallel for each vertex.
@@ 6.4.1 Triangle counting in the host graph
6.4.2 Counting subgraph of radius one
We begin by explaining the procedure for counting the number of subgraphs having a dominating vertex (radius one). For this purpose, we fix a dominating vertex $u$. If a subgraph exists, then the dominating vertex must be mapped to some vertex. We iteratively map the dominating vertex to each vertex in the host graph and count the number of patterns in the neighbourhood of the dominating vertex.
Here, we present an algorithm for counting patterns of radius one 6. Note that running $k$-WL on the entire graph takes $O(k^{2}\cdot n^{k+1}\log n)$ time and $O(n^{k})$ space, whereas when we run locally, it requires less time and space. Suppose we run only on the neighbourhood of each vertex. Then, it requires $\sum_{v\in V(G)}(deg(v))^{k+1}\log(deg(v))$ and space $O(max_{i}(deg(v_{i}))^{k}+n).$ More specifically, suppose the given graph is $r$-regular. Then it requires $O(r^{k+1}\log(r)n)$ time and $O(r^{k}+n)$ space. Therefore, if the graph is sparse, then we can implement Local $k$-WL for a larger value of $k$. We can see that running $k$-WL locally is not dependent on the size of the graph exponentially. However, it is not feasible to run $k$-WL on the entire graph for a larger value of $k.$
@@ 6.4.2 Counting subgraph of radius one
6.4.3 Counting subgraphs of radius r
Here, in 7, we describe how to count the number of subgraphs of radius $r.$ We iterate over all the vertices and take the $r$-hop neighbourhood around that vertex, say $v,$ and choose a suitable $k$ according to the structure of the pattern that can guarantee the count of subgraphs in the local substructure.
@@ 6.4.3 Counting subgraphs of radius r
6:   if #\# vertices at layer $i$ from $v$ << #\# vertices at layer $i$ from $u$ then
@@
7 Methodology
This section describes our unified subgraph counting strategy, combining localized Weisfeiler-Leman (WL) algorithms, differentiable learning, and a novel fragmentation approach. The framework is broken into three primary components: (A) Pattern Learning, (B) Local Count Learning, and (C) Global Count Learning, as illustrated in Figure 2.
Figure 2: Overview of the framework. (A) Pattern Learning: Each node’s $r$-hop neighbourhood, with directed edges marking the root, is input to a model that predicts relevant subpatterns. (B) Local Count Learning: Using predicted patterns, the model estimates subpattern counts per root node. (C) Global Count Learning: The total count is normalized to adjust for possible overcounting.
7.1 Pattern Learning
For a given subgraph pattern $P$ (e.g., triangle, $3$-star, or tailed triangle), the local subgraph $G_{v}^{r}$ centered at node $v$ must be preprocessed to match $P$’s structure. We augment $G_{v}^{r}$ by directing edges outward from $v$ to clearly indicate the root. Modified $G_{v}^{r}$ is passed to a GNN, which produces node and edge embeddings. These embeddings are used to classify edges as included/excluded based on their relevance to the target pattern, yielding a pruned subgraph suited as input to the next component.
7.2 Local Count Learning
Each root node’s updated local subgraph (from Pattern Learning) is used to estimate the "local count"—the number of occurrences of a subpattern in $G_{v}^{r}$. The local counting rule depends on the motif: for triangles, count edges among neighbours; for $k$-stars, compute $\binom{d_{v}}{k}$ where $d_{v}$ is node $v$’s degree. For arbitrary patterns, ground truth local counts are generated as supervision. The model is trained to predict these counts using a GNN applied to the modified $G_{v}^{r}$.
7.3 Global Count Learning
Summing local counts across all roots may overcount global subgraph occurrences due to overlaps. Hence, a final normalization step is applied. We use a learnable normalization function to infer the correction factor needed to yield the true global count, enabling an end-to-end differentiable pipeline for subgraph counting.
7.4 Fragmentation
For complex patterns beyond primitives like triangles or $k$-stars, we introduce the fragmentation technique. Instead of directly counting $P$, which may be computationally intensive, we recursively decompose $P$ into simpler subpatterns ("fragments") whose counts can be efficiently determined. The total count of $P$ is obtained by suitably combining the counts of these fragments, often requiring only small values of $k$ in the local $k$-WL hierarchy.
For example, the tailed triangle can be fragmented into a pendant vertex (the key/root) and an attributed triangle, as shown in Figure 3. For chordal $C_{4}$ cycles, the pattern is broken into $2$-stars and triangles within the neighbourhood of a key vertex, adjusting counts to avoid overcounting $K_{4}$s. The approach applies for any pattern whose complexity can be reduced through such decomposition.
Figure 3: Fragmentation example for a tailed triangle: identifying fragments (isolated node, edge) in the subgraph surrounding a root node.
Fragmentation Algorithm:
Given a host graph $G$ and a pattern $P$ to be counted, select a key vertex in $P$ and set the radius $r$. For each node $v\in V(G)$, extract the $r$-hop neighbourhood $G_{v}^{r}$. Decompose $P$ into subpatterns $\{P_{1},...,P_{l}\}$, each counted by a learned model $M_{i}$. Store local counts in arrays, aggregate via learnable combination functions $\alpha$, $\beta$ (for subpattern and root aggregation, respectively), and finally apply a normalization function $\gamma$ to compute the global count.
@@ 7.4 Fragmentation
1:$G$; $\mathbb{P}$: List of subpatterns; $M^{\text{pattern}}_{i}$: learned model for generating pattern corresponding to $P_{i}\in\mathbb{P}$; $M^{\text{count}}_{i}$: learned model for counting $P_{i}\in\mathbb{P}$;
[…]
3:for each node $v\in V(G)$ do
4:  $H_{v}=G_{v}^{r}\setminus\{v\}$
5:  $b\leftarrow[]$
6:  for each pattern $P_{i}\in\mathbb{P}$ do
7:   $c\leftarrow[]$
8:   for each node $u\in H_{v}$ do
9:     $\mathcal{P}_{u}=M^{\text{pattern}}_{i}(H_{v},u)$
10:     $c.append(M^{\text{count}}_{i}(\mathcal{P}_{u}))$
11:   end for
12:   $b.append(\alpha(c))\qquad\#$Learnable function
@@
14:  $a.append(\beta(b))\qquad\#$Learnable function
@@
[…]
17:return CountCount
@@
Using the fragmentation method, we can count induced tailed triangle, chordal $C_{4}$ and 3star,3-star, by running Local $1$-WL.
@@
For tailed triangle, we fix the pendant vertex as the key vertex (refer to 3). Now, we have to look for the count of triangles such that exactly one of the node of triangle is adjacent to the key vertex. We find induced subgraph of 22-hop neighborhood for each key vertex. Now, we color the vertices at distance $i$ by color $i$. Then, the problem of counting the number of tailed triangles reduces to counting the number of colored triangles in the colored graph such that one node of triangle is colored $1$ and remaining two nodes are colored $2$. We can find the count of colored triangles using 11-WL on the induced subgraph by Corollary 1. This number of count of colored triangle is same as $IndCount_{(u,v}(tailed-triangle,G_{v}^{r})$. Now, using 3, we can say that fragmentation technique can count tailed triangles appearing as induced subgraphs using 11-WL.
Consider the pattern, chordal $C_{4}.$ We have proved that 11-WL can count the number of subgraphs of chordal $C_{4}.$ So, to count the number of induced subgraphs of chordal $C_{4},$ we only have to eliminate the count of $K_{4}.$ When we fix one vertex of degree $3$ as key vertex, we can easily compute the count of $K_{1,2}$ in the neighborhood. Now, we have to eliminate all three tuples appearing as triangles in the neighborhood of the key vertex. We can easily count the number of triangles in the neighborhood of each vertex. This gives the exact count of chordal $C_{4}$ appearing as subgraph in the local structure. Using 3, we can find $IndCount(ChordalC_{4},G)$.
Consider the pattern, 3star3-star. Here, we choose any of the pendant vertex as key vertex. Now, we have to compute $K_{1,2},$ where one center vertex of the star is connected to the key vertex. We can easily count the number of colored $K_{1,2}$ in 22-hop neighborhood of the key vertex. However, a triangle can be also be included in this count. So, we have to eliminate the $3$ tuples forming a triangle. Again, using the approach discussed above, we can count the number of colored triangles and this will output the exact count of colored induced $K_{1,2}$. Again, using lemma 3, we can find IndCount(3star,G).IndCount(3-star,G).
Table 2: Fragmentation for patterns for size of at most 4
@@
G3 [Uncaptioned image] $\frac{\sum_{v}IndCount_{(u,v)}(G_{1},G-G[N_{G}[v]])}{3}$
@@
Theorem 11.
Proof.
In 2, we describe how the fragmentation technique can be leveraged to count all the induced subgraphs of size $4$. This shows that for $k=1$, the fragmentation technique is more expressive than $S_{k+3}$. Also, we can enlist more graphs where the larger pattern can be seen as the union of fragments of smaller patterns. Using this, we can see that it covers all the graphs that were mentioned in (chen2020can). One can see that all the formulae use the function that can be computed by 11-WL. The number of vertices and the number of edges can easily be computed after doing one iteration of 11-WL. Also, the degree sequence can be computed after running $1$-WL for one iteration. All other functions(formulae) are just the linear combination of the functions computed earlier, the number of vertices or the number of edges. ∎
Expressiveness Guarantee:
Using fragmentation, we show that all patterns of up to $4$ nodes can be exactly counted using only Local $1$-WL. For specific motifs (tailed triangle, chordal $C_{4}$, $3$-star), the decomposition ensures algorithmic tractability and high expressiveness (see Theorems 10 and 11).
The fragmentation technique thus enables scalable, modular, and highly expressive subgraph counting, leveraging learnable models for each subcomponent while ensuring practical feasibility and end-to-end differentiable training.
@@ 8 Theoretical Comparison of Graph Neural Networks
n this section we give a comparison of the time and expressiveness between the GNN hierarchies proposed in (papp2022theory) and our methods. From 3, it is clear that $k$-WL is less expressive than Local $k$-WL. Also, we have shown that the space and time required by Local $k$-WL are less compared to $k$-WL.
@@
$S_{k}$ is a model in which a vertex is given an attribute based on the number of connected induced subgraphs of size at most $k,$ the key vertex. Even though it searches locally, the number of non-isomorphic graphs may be too many for a small value of $k$. Suppose the radius of the induced subgraph is $r$; then it has to search up to the $r$-hop neighborhood. Using brute force would require $O(n_{1}^{k})$-time to count the number of induced subgraphs of size $k,$ for each individual induced subgraph. To improve the time complexity, it either stores the previous computation, which requires a lot of space or further recomputes from scratch. Thus, it would require $O(t_{k}\times n_{1}^{k})$, where $t_{k}$ is the number of distinct graphs upto isomorphism of at most $k$-vertices. Using 6, one can easily see that running $k$-WL locally is more expressive than $S_{k+2}.$
@@
The $N_{k}$ model has a preprocessing step in which it takes the $k$-hop neighborhood around vertex $v,$ and gives attributes based on the isomorphism type. In a dense graph, it is not feasible to solve the isomorphism problem in general, as the size of the induced subgraph may be some function of $n$. Also, using the best-known algorithm for graph isomorphism by babai2016graph, the time required is $O(n_{1}^{O(\log n_{1})}).$ However, running Local $k$-WL would require $O(n_{1}^{k})$. Also, there are rare examples of graphs that are 33-WL equivalent and non-isomorphic. So, if we run $3$-WL locally, then most of times expressive power matches with $N_{k}.$
@@
The $M_{k}$ model deletes a vertex $v$ and then runs 1WL.1-WL. papp2022theory proposed that instead of deleting the vertices, $k$ vertices are marked in the local neighborhood and showed that it is more expressive than deletion. It identifies $k$ set of vertices in the local $r$-hop neighborhood of the graph. It would require $O(n_{1}^{(k+2)}\log(n_{1}))$ time as it has $O(n_{1}^{k})$ many possibilities of choosing $k$ many vertices. It requires $O(n^{2}\log n)$ time to run 11-WL. The same time is required for Local $(k+1)$-WL. Also, it is known that with any $l$ markings and running $k$-WL is less expressive than running $(k+l)$-WL on the graph (furer2017combinatorial). So, if we plug in the value, we can see that running Local $k$-WL is more expressive than doing $l$ marking and running 11-WL. One can get an intuition by comparing with the $(k+1)$ bijective pebble game. If we mark the vertices, then the marking pebbles get fixed and give less power the to spoiler. However, just running $k$-WL the spoiler is free to move all the pebbles. We present a simple proof that Local $k$-WL is at least as expressive as $M_{k-1}$.
@@
Let $G_{v}^{r}$ and $G_{u}^{r}$ be the induced subgraphs around the $r$-hop neighborhood for vertices $v$ and $u$, respectively. Let $M_{k}$ distinguish $G_{v}^{r}$ and $G_{u}^{r}$. We claim that Local $k$-WL can also distinguish the graph $G_{v}^{r}$ and $G_{u}^{r}$. To prove our claim, we use the pebble bijective game. $G_{v}^{r}$ is distinguished because there exists a tuple $(v_{1},v_{2},....v_{k-1})$ such that marking these vertices and running $1$-WL on the graph gives stabilised coloring of the vertices that does not match with that of $G_{u}^{r}.$ Now, consider two games. One game corresponds to 11-WL and another to Local $k$-WL. For the first (k1)(k-1) moves, the Spoiler chooses to play and places pebbles at $(v_{1},v_{2},....v_{k-1}).$ After that, in both games, there are two pebbles and the position of both games are the same. Let $S_{1}$ and $D_{1}$ be the spoiler and Duplicator in the $(k+1)$ bijective pebble game, and $S_{2}$ and $D_{2}$ be the spoiler and Duplicator in $2$ bijective pebble game. $S_{1}$ will follow the strategy of $S_{2}$ and $D_{2}$ follows the strategy of $D_{1}.$ We prove by induction on the number of rounds. Our induction hypothesis is that the position of games in both the games is the same and if $S_{2}$ wins, then $S_{1}$ also wins.
@@
9 Experiments